OpenStudy (pokemon23):

solve for b in terms of a and L: =a+b show all work step by step show all work explain to me how you got the answer free medal solve for b in terms of a and L: =a+b show all work step by step show all work explain to me how you got the answer free medal @Mathematics

OpenStudy (anonymous):

okay lets go through this together. i need you to answer my questions. L = a+b correct?

OpenStudy (anonymous):

subtract a from both sides and tell me what you get.

OpenStudy (pokemon23):

ok

OpenStudy (anonymous):

go on. write the expression you get when you subtract a from both sides

OpenStudy (pokemon23):

L-a=b

OpenStudy (anonymous):

right so L-a = b in other words L minus a equals b in yet other words, b equals L minus a. therefore b = L-a

OpenStudy (pokemon23):

sorry im retarded in math

OpenStudy (pokemon23):

I'm actually correct

OpenStudy (pokemon23):

what is the next step

OpenStudy (anonymous):

and voila! you have an expression for b in terms of L and a. in other words, you have successfully managed to express b using only L and a. so if you have an expression like 15b, you can simply write it as 15(L-a) because b = L-a.

OpenStudy (anonymous):

that is all there is to it.

OpenStudy (pokemon23):

I see can you help me with the other questions

OpenStudy (pokemon23):

I like how you explain things

OpenStudy (pokemon23):

Make math much easier for me

OpenStudy (anonymous):

post then on the left. I will get to them

OpenStudy (pokemon23):

ok thanks